OpenStudy (anonymous):

how to find the 15th term for the following sequence : 15,8,1,-6,-13

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

There is a sequence of subtracting 7 every time so you would keep going until you get to the fifteenth term

OpenStudy (anonymous):

15, 8, 1, -6, -13, -20, -27, -34, -41, -48, -55,-62, -69, -76, -83

OpenStudy (anonymous):

-83 is the fifteenth term

OpenStudy (texaschic101):

first we need to find common difference between the numbers...to do that we subtract the first term from the second term....8 - 15 = -7 you will find that the common difference between all these numbers is -7 now to finish this, we have a formula an = a1 + (n - 1) d a1 = first term = 15 d = common difference = -7 n = number of term you want to find = 15 now we sub a15 = 15 + (15 - 1)(-7) a15 = 15 + (14) * -7 a15 = 15 + (-98) a15 = 15 - 98 a15 = - 83 so the 15th term is -83